# Prospects of constraining on the polarizations of gravitational waves from binary black holes using space- and ground-based detectors

### v1

@@ III.3 Response function
Figure 5: Angular response functions for the combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ at LISA detector coordinates $(\phi_{d},\theta_{d})$. The left panel shows the response functions for GWs at 1 mHz, where the breathing and longitudinal modes are almost identical. The right panel displays the response at twice the transfer frequency, $2f_{*}=38\ \mathrm{mHz}$, where the degeneracy between the breathing and longitudinal modes is broken at higher frequencies.
A dual-arm Michelson interferometer detects GWs by measuring the relative change in the length of its two arms. We describe the detector’s response to GWs using the method outlined in Refs. [45, 27]. As shown in Fig. 4, the detector coordinates are constructed using orthogonal unit vectors $\hat{x},\hat{y},\hat{z}$}, and the GW coordinates are constructed using $\hat{p},\hat{q},\hat{w}$}. Here, $\hat{w}$ represents the propagation direction of the GW, and the unit vector $\hat{\Omega}=-\hat{w}$ represents the position of the GW source. In a triangular detector, the angle $\gamma$ between the two arms is $60^{\circ}$, while in an L-shaped detector, it is $90^{\circ}$. For GWs, an additional rotational degree of freedom can be fixed by specifying the polarization angle $\psi$, ultimately using $\hat{m},\hat{n},\hat{w}$} to describe the GW. The angular response function $F^{A}$ in Eq. (1) can be expressed as
@@
$e_{ij}^{B}$ $=\hat{m}_{i}\hat{m}_{j}+\hat{n}_{i}\hat{n}_{j},$ $e_{ij}^{L}$ $=\hat{w}_{i}\hat{w}_{j}.$
And the detector tensor $D^{ij}$ can be writen as [60]
@@
$\mathcal{T}(f,\hat{a}\cdot\hat{b})=$ $\frac{1}{2}\left\{\sin c\left[\frac{f}{2f_{*}}(1-\hat{a}\cdot\hat{b})\right]\right]$ (9)
@@
where $\operatorname{sinc}(x)=\sin x/x$, $f_{*}=c/(2\pi L)$ is the transfer frequency and $L$ is the arm length of the detector. The transfer frequencies for space-based detectors are 19 mHz for LISA, 16 mHz for Taiji, and 275 mHz for TianQin. Due to the arm length, the transfer frequency of ground-based detectors is significantly higher than their sensitivity frequency band.
The angular response functions $F^{A}(\lambda,\beta,\psi)$ or $F^{A}(\alpha,\delta,\psi)$ are obtained by substituting the detector and ecliptic/equatorial coordinates into Eqs. (6)-(9) using Euler rotation conversion. To present the angular response function concisely, we introduce the combined tensor and vector modes:
@@
According to Ref. [27], the angular response functions of combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ are independent of the polarization angle $\psi$. In this section, we study the response of different modes using these four independent modes of $\psi$ and the six basic modes for subsequent research and calculations. On this basis, we calculate the angular response functions for two different frequencies in the LISA detector coordinates to plot Fig. 5.
According to Eq. (9), at the low-frequency limit $f\ll f_{*}$, $\mathcal{T}\to 1$ leads to $F^{B}=-F^{L}$ [27]. Ground-based detectors operate within sensitivity frequency bands below the low-frequency limit, making it impossible to distinguish between these two modes. Moreover, space-based detectors may break through the low-frequency limit and resolve this degeneracy. Figure. 5 shows that at the low-frequency limit, the angular response functions for the breathing mode and longitudinal mode are degenerate at every position. Beyond the transfer frequency, the optimal response position of the breathing mode shifts from $\phi_{d}=45^{\circ},135^{\circ},225^{\circ},315^{\circ}$ to $\phi_{d}=90^{\circ},270^{\circ}$. This shift, which differs from the longitudinal mode’s optimal response position, makes it possible to break the degeneracy of these two modes.
@@
Figure 5 illustrates the response at two specific frequencies. To further examine the relationship between response and frequency, we introduce the averaged angular response function averaged over the source locations [27]:
@@
where $A=\{T,V,B,L\}$ denotes four modes independent of $\psi$. We also introduce effective strain noise to measure the detector’s sensitivity to different modes varying with frequency:
@@
Figure 6: The averaged response functions $R_{A}(f)$ and effective strain noise $h^{A}_{eff}(f)$ of LISA’s four modes vary with frequency. We calculate $R_{A}(f)$ and $h^{A}_{eff}(f)$ for each frequency point through integration. The purple vertical dashed line represents LISA’s transfer frequency $f_{*}=19.1\ \mathrm{mHz}$, and the black dashed line is a fitting approximation of $R_{A}(f)$ (cf. Ref. [27]).
After our calculations, we derive the averaged response functions and effective strain noise for LISA’s four modes, as shown in Fig. 6. At the low-frequency limit, the averaged response function remains constant, with $R_{T}=R_{V}=2(\sin^{2}\gamma)/5$ and $R_{B}=R_{L}=(\sin^{2}\gamma)/15$. As the transfer frequency approaches, $R_{A}$ starts to decrease, showing three distinct damping trends. $R_{T}$ and $R_{V}$ diverge with increasing frequency, similar to $R_{B}$ and $R_{L}$. Additionally, $R_{T}$ and $R_{B}$ show differences at the low-frequency limit. As the frequency increases, the values of those two modes tend to converge. Note that while the average value of these two modes is the same, their responses at individual positions differ, preventing degeneracy between $R_{T}$ and $R_{B}$.
Based on Ref. [27], we analyze the detector’s capability to detect various polarizations using the response function. For ground-based detectors, the sensitive frequency band is below the transfer frequency, preventing the distinction between the breathing mode and the longitudinal mode. Space-based detectors, however, can break this degeneracy at non-low-frequency limits. In Sec. V, we simulate GW signals to evaluate the performance of ground- and space-based detectors in constraining polarization.
@@ IV.1 Data analysis
Figure 7: Multiband observation of BBHs with ground- and space-based detectors. The vertical axis represents the dimensionless characteristic strain $\sqrt{fS_{n}(f)}$. The BBH curve is derive from Ref. [53] and is intended for illustration purposes only, not for calculation purposes. For luminosity distance selection, we set SBBH at $z=0.01\ (D_{L}=44.6\ \mathrm{Mpc})$ and MBHB at $z=1\ (D_{L}=6.8\ \mathrm{Gpc})$. The cyan line represents the space detection source, while the olive line represents the ground detection source. The three vertical dashed lines represent the transfer frequencies of the three space-based detectors.
@@
where $\tilde{a}(f)$ and $\tilde{b}(f)$ are the Fourier transforms of $a(t)$ and $b(t)$ respectively. In our study, we consider fourteen GW parameters in total, including ppE parameters, which are
$\bm{\xi}=\{t_{c},m_{1},m_{2},z,\iota,\Phi_{0},\phi_{e},\theta_{e},$ (15)
@@
where $t_{c}$ is the coalescence time, $z$ is the redshift of the source, $(\phi_{e},\theta_{e})$ is the sky position, representing ecliptic $(\lambda,\beta)$ or equatorial $(\alpha,\delta)$ coordinates.
For assessing the limitations of the detector on different polarizations and the uncertainty in estimating all parameters, we use the FIM method, defined as
@@
When calculating the FIM in Eq. (16), we use the numerical differentiation approximation from Refs. [44, 63]. Additionally, for the observation network, the total SNR and FIM are obtained by summing the inner products calculated by each detector.
@@ IV.2 BBH source selection
Space- and ground-based detectors have different sensitive frequency bands, resulting in the detection of various GW sources. For BBH sources, ground-based detectors primarily observe SBBHs capturing the complete CBC process of three phases, whereas space-based detectors mainly observe that of MBHBs. Additionally, space-based detectors can detect SBBHs, as many SBBHs inspiral in the low-frequency band before merger, entering the sensitivity band of these detectors. The different BBH sources we selected are presented in Fig. 7.
For all CBC processes, since additional polarizations contribute more significantly in the inspiral phase and Eq. (2) does not apply to the merger and ringdown phases, our waveform model concentrates on the inspiral phase before the binary reaches the innermost stable circular orbit (ISCO). The frequency of ISCO is given by [36]
@@
Ground-based detectors typically observe SBBH GWs entering the frequency band a few seconds to minutes before merger. Thus, we select three typical SBBHs with equal mass ratios and total masses of $3\ \mathrm{M_{\odot}}$, $20\ \mathrm{M_{\odot}}$, and $100\ \mathrm{M_{\odot}}$, respectively. We calculate the waveforms for the 10 minutes leading up to ISCO. For MBHB sources observed by space-based detectors, we select MBHBs with three typical masses of $10^{5}\ \mathrm{M_{\odot}}$, $10^{6}\ \mathrm{M_{\odot}}$, and $10^{7}\ \mathrm{M_{\odot}}$, calculating waveforms for the 3 months before reaching ISCO. For multiband observations, we extend two types of SBBHs with masses of $20\ \mathrm{M_{\odot}}$ and $100\ \mathrm{M_{\odot}}$ to the low-frequency band. Those waveforms last for a year before reaching 0.1 Hz. After 0.1 Hz, those SBBHs would take 3 months ($20\ \mathrm{M_{\odot}}$) and 6 days ($100\ \mathrm{M_{\odot}}$) to enter the ground-based detector observation frequency band, providing an opportunity to study multiband observations.
@@
For the selection of sky positions for BBH sources, we use identical parameters for each source, generating 100 distributions consistent with Table 1. To study the influence of inclination angle, we consider 45 different inclination angles, resulting in 4500 calculations of SNR and FIM for each BBH source, with the results presented in Secs. V.1 and V.2. Finally, for selecting fiducial values for the ppE parameter, we set it to $10^{-4}$ (cf. [36]).
@@ V.1 Results with inclination
We specifically examine the relationship between the uncertainty of ppE parameters and the inclination angle. According to Eq. (2), we can define the distribution of inclination of the tensor mode $p_{T}(\iota)$ as
@@
Similarly, we can determine that the distributions of vector mode $p_{V}(\iota)$ and scalar mode $p_{S}(\iota)$ are
@@
We present the distributions of those three modes in Fig. 8.
@@
Figure 9: Relationship between SNR and parameter uncertainty with inclination angle variation. The vertical axis is represented in logarithmic scale. SNR is calculated based on the GR case, and each line in the graph represents the median value. The results for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are identical, so only one is shown. The left panel includes LISA (solid line) and TianQin (dashdot line), while the right panel includes LIGO (solid line) and ET (dashdot line).
@@
From Fig. 8, it is evident that the distribution of scalar mode with inclination is opposite to that of tensor and vector modes, reaching its maximum value at $\iota=\pi/2$ and minimum at $\iota=0,\pi$. From the perspective of inclination angle influence alone, the scalar mode is superior to the other two modes at $\iota=\pi/2$, while tending towards 0 at $\iota=0,\pi$. Tensor and vector modes exhibit the same trend of change, with tensor modes showing a greater amplitude of variation. Furthermore, we calculate the SNR and FIM to obtain the results for the uncertainty of the ppE parameters, as shown in Fig. 9.
@@
As the distribution in Fig. 8, the variation of the scalar mode with inclination in Fig. 9 is opposite to other modes. Furthermore, due to the distribution of scalar mode being zero at $\iota=0,\pi$, there is almost no signal near those two values, resulting in the parameter uncertainty of the scalar mode in Fig. 9 approaching infinity. In addition, considering the low-frequency limit, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of a $10^{5}\ \mathrm{M_{\odot}}$ MBHB with low-SNR are smaller than that of a $10^{7}\ \mathrm{M_{\odot}}$ MBHB with high-SNR. Moreover, due to the superior sensitivity of LISA compared to TianQin and the lower transfer frequency of LISA, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of LISA are significantly better than that of TianQin. For the degeneracy caused by arm length, ground-based detectors cannot distinguish between the breathing mode and the longitudinal mode, so the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are much greater than that of space-based detectors and becomes invalid.
@@
Besides the scalar mode, a higher SNR signifies a more robust signal, typically leading to reduced parameter uncertainties. In our calculations, SNR is calculated based on the GR case, where the contribution of tensor modes outweighs that of other modes. Thus, the SNR value is primarily influenced by the tensor modes. In contrast to tensor and scalar modes, the uncertainty of the vector mode is notably higher at $\iota=\pi/2$ in TianQin. This is because TianQin’s fixed orientation restricts the response function’s variability at that specific angle, leading to a pronounced increase in $\Delta\alpha_{V}$ near $\iota=\pi/2$.
@@
The transfer frequency not only directly impacts the degeneracy between breathing and longitudinal modes but also affects other aspects. The transfer function $\mathcal{T}$ is directly related to frequency, providing constraints beyond the low-frequency limit. $\alpha_{Q}$ and $\alpha_{D}$ determine the variation of the orbital angular frequency, and the influence of $\mathcal{T}$ can reduce parameter uncertainty. Therefore, the $\Delta\alpha_{Q}$ of $10^{5}\ \mathrm{M_{\odot}}$ MBHB is smaller than that for a $10^{6}\ \mathrm{M_{\odot}}$ MBHB, and $\Delta\alpha_{D}$ for both MBHBs also varies with the relationship near $\iota=\pi/2$. Because of the low-frequency limit, ground-based detectors are not affected by $\mathcal{T}$, causing the uncertainties $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ for SBBHs to depend entirely on SNR. This section focuses on presenting and analyzing the relationship between the uncertainty of ppE parameters and the inclination angle for two typical space-based and ground-based detectors.

### v2

@@ III.3 Response function
Figure 5: Angular response functions for the combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ at LISA detector coordinates $(\phi_{d},\theta_{d})$. The left panel shows the response functions for GWs at 1 mHz, where the breathing and longitudinal modes exhibit almost indistinguishable characteristics. The right panel displays the response at twice the transfer frequency, $2f_{*}=38\ \mathrm{mHz}$, where a distinction between the breathing and longitudinal modes emerges as frequencies increase.
The detection of GWs by a dual-arm Michelson interferometer is achieved by measuring the relative change in length between its two arms. We describe the detector’s response using the method outlined in Refs. [48, 27]. The detector coordinates, as illustrated in Fig. 4, are constructed using orthogonal unit vectors $\hat{x},\hat{y},\hat{z}$}. Similarly, the GW coordinates are constructed using $\hat{p},\hat{q},\hat{w}$}. Here, $\hat{w}$ represents the propagation direction of the GW, while the unit vector $\hat{\Omega}=-\hat{w}$ denotes the position of the GW source. The angle $\gamma$ between the two arms is $60^{\circ}$ in a triangular detector and $90^{\circ}$ in an L-shaped detector. For GWs, an additional rotational degree of freedom can be fixed by specifying the polarization angle $\psi$, ultimately using $\hat{m},\hat{n},\hat{w}$} to describe the GW. The angular response function $F^{A}$ in Eq. (1) can be expressed as
@@
$e_{ij}^{B}$ $=\hat{m}_{i}\hat{m}_{j}+\hat{n}_{i}\hat{n}_{j},$ $e_{ij}^{L}$ $=\hat{w}_{i}\hat{w}_{j},$
and the detector tensor $D^{ij}$ can be written as [67]
@@
$\mathcal{T}(f,\hat{a}\cdot\hat{b})=$ $\frac{1}{2}\biggr{\{}\operatorname{sinc}\left[\frac{f}{2f_{*}}(1-\hat{a}\cdot\hat{b})\right]$ (9)
@@
where $\operatorname{sinc}(x)=\sin x/x$, $f_{*}=c/(2\pi L)$ is the transfer frequency and $L$ is the arm length of the detector. The transfer frequencies for space-based detectors are 19 mHz for LISA, 16 mHz for Taiji, and 275 mHz for TianQin. Due to the arm length, the transfer frequency of ground-based detectors exceeds their sensitivity frequency band significantly.
The angular response functions $F^{A}(\lambda,\beta,\psi)$ or $F^{A}(\alpha,\delta,\psi)$ are derived by substituting the detector and ecliptic/equatorial coordinates into Eqs. (6$\sim$(9) using Euler rotation conversion. To present the angular response function concisely, we introduce the combined tensor and vector modes:
@@
The angular response functions of the combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ are independent of polarization angle $\psi$ [27]. The response of different modes is studied using these four independent modes of $\psi$ and the six basic modes for subsequent research and calculations. On this basis, we calculate the angular response functions for two different frequencies in the LISA detector coordinates to generate Fig. 5.
According to Eq. (9), in the low-frequency regime $f\ll f_{*}$, the limit $\mathcal{T}\to 1$ implies $F^{B}=-F^{L}$ [27]. The sensitivity frequency bands of ground-based detectors are below this low-frequency limit, making it impossible to distinguish between these two modes. The potential of space-based detectors to surpass the low-frequency limit and resolve this degeneracy is significant. The angular response functions for the breathing mode and longitudinal mode are degenerate at every position in Fig. 5, indicating their degeneracy at the low-frequency limit. Beyond the transfer frequency, the optimal response position of the breathing mode shifts from $\phi_{d}=45^{\circ},135^{\circ},225^{\circ},315^{\circ}$ to $\phi_{d}=90^{\circ},270^{\circ}$. The shift, which is distinct from the optimal response position of the longitudinal mode, allows for the differentiation between these two modes.
@@
The response at two specific frequencies is illustrated in Fig. 5. In order to further investigate the relationship between response and frequency, we introduce the angular response function averaged over the source locations [27],
@@
where $A=\{T,V,B,L\}$ denotes four modes independent of $\psi$. The detector’s sensitivity to different frequency modes can be measured by introducing effective strain noise,
@@
Figure 6: The averaged response functions $R_{A}(f)$ and effective strain noise $h^{A}_{eff}(f)$ of LISA’s four modes vary with frequency. We calculate $R_{A}(f)$ and $h^{A}_{eff}(f)$ for each frequency point through integration. The purple vertical-dashed line represents LISA’s transfer frequency $f_{*}=19.1\ \mathrm{mHz}$, and the black dashed line is a fitting approximation of $R_{A}(f)$ (cf. Ref. [27]).
Through the calculations, we obtain the averaged response functions and effective strain noise for LISA’s four modes, as shown in Fig. 6. At the low-frequency limit, the averaged response function remains constant, with $R_{T}=R_{V}=2(\sin^{2}\gamma)/5$ and $R_{B}=R_{L}=(\sin^{2}\gamma)/15$. The decrease in $R_{A}$ becomes evident as the transfer frequency increases, exhibiting three distinct damping trends. $R_{T}$ and $R_{V}$ diverge as the frequency rises, similar to the behavior of $R_{B}$ and $R_{L}$. The differences between $R_{T}$ and $R_{B}$ become apparent at low frequencies, but as the frequency increases, these two modes tend to converge. The average values of these two modes are identical, but their responses at individual positions differ, preventing degeneracy between $R_{T}$ and $R_{B}$.
Based on Ref. [27], we analyze the detector’s capability to detect various polarizations using the response function. The sensitive frequency band of ground-based detectors is below the transfer frequency, making it difficult to distinguish between the breathing mode and longitudinal mode. The degeneracy at nonlow-frequency limits can be resolved by space-based detectors. In Sec. V, we simulate GW signals to assess the capability of ground- and space-based detectors in polarization constraint.
@@ IV.1 Data analysis
Figure 7: Multiband observation of BBHs with ground- and space-based detectors. The vertical axis represents the dimensionless characteristic strain $\sqrt{fS_{n}(f)}$. The BBH curve is derive from Ref. [60] and is intended for illustration purposes only, not for calculation purposes. For luminosity distance selection, we set SBBH at $D_{L}=44.6$ Mpc ($z=0.01$ in $\mathrm{\Lambda CDM}$ cosmological model) and MBHB at $D_{L}=6.79$ Gpc ($z=1$). The cyan line represents the space detection source, while the olive line represents the ground detection source. The three vertical dashed lines represent the transfer frequencies of the three space-based detectors.
@@
where $\tilde{a}(f)$ and $\tilde{b}(f)$ are the Fourier transforms of $a(t)$ and $b(t)$, respectively. The total number of GW parameters considered in our study is fourteen, including ppE parameters, which are
$\bm{\xi}=\{t_{c},m_{1},m_{2},D_{L},\iota,\Phi_{0},\phi_{e},\theta_{e},$ (15)
@@
where $t_{c}$ is the coalescence time, $D_{L}$ is the luminosity distance of the source, $(\phi_{e},\theta_{e})$ is the sky position, representing ecliptic $(\lambda,\beta)$ or equatorial $(\alpha,\delta)$ coordinates.
To evaluate the detectors’ performance on different polarizations, we use the FIM:
@@
When calculating the FIM, we use the numerical differentiation approximation as described in Refs. [47, 70]. The total SNR and FIM for the observation network are obtained by summing the inner products calculated by each detector.
@@ IV.2 BBH source selection
The distinct sensitive frequency bands of space- and ground-based detectors facilitate the detection of various GW sources. Ground-based detectors can observe SBBH’s whole CBC in three phases, while space detectors do the same for MBHB. Additionally, space detectors can also observe the inspiraling SBBH in the low-frequency band. The selected BBH sources are shown in Fig. 7.
The proposed waveform model focuses on the inspiral phase before the binary reaches the innermost stable circular orbit (ISCO). Additional polarizations plays a more significant role in this phase than in merger and ringdown phases. The frequency of ISCO is given by [36]
@@
Ground-based detectors typically observe the entry of SBBH GWs into the frequency band a few seconds to minutes prior to their merger. Therefore, we have selected three typical SBBHs with equal mass ratios and total masses of $3\ \mathrm{M_{\odot}}$, $20\ \mathrm{M_{\odot}}$, and $100\ \mathrm{M_{\odot}}$, respectively. The waveforms for the 10 minutes preceding ISCO are computed. For MBHB sources observed by space-based detectors, we select MBHBs with three representative masses: $10^{5}\ \mathrm{M_{\odot}}$, $10^{6}\ \mathrm{M_{\odot}}$, and $10^{7}\ \mathrm{M_{\odot}}$. The waveforms are calculated for the three months before ISCO. For multiband observations, the SBBHs with masses of $20\ \mathrm{M_{\odot}}$ and $100\ \mathrm{M_{\odot}}$ are extended to the low-frequency band, lasting for a year before reaching 0.1 Hz. The SBBHs beyond this threshold would take approximately three months ($20\ \mathrm{M_{\odot}}$) and six days ($100\ \mathrm{M_{\odot}}$) to enter the frequency band of ground-based detectors, providing an opportunity for conducting multiband observations.
@@
We generate 100 BBH sources. The sky positions for BBH sources are selected with identical parameters, following the distribution in Table 1. In order to investigate the impact of inclination angle, we consider 45 angles, resulting in a total of 4500 calculations of SNR and FIM for each BBH source. The corresponding results are presented in Secs. V.1 and V.2. Finally, we set the fiducial value for the ppE parameter to $10^{-4}$ (cf. [36]).
@@ V.1 Results with inclination
The relationship between ppE parameter uncertainty and the inclination angle is specifically examined. By referring to Eq. (2), we can define the distribution of inclination for the tensor mode as
@@
The distributions of vector mode $p_{V}(\iota)$ and scalar mode $p_{S}(\iota)$ can be determined similarly,
@@
The distributions of these three modes are presented in Fig. 8.
@@
Figure 9: Relationship between SNR and parameter uncertainty with inclination angle variation. The vertical axis is represented in logarithmic scale. The SNR is calculated based on the GR case, and each line in the graph represents the median value. The results for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are identical, therefore only one is shown. The left panel includes LISA (solid line) and TianQin (dashdot line), while the right panel includes LIGO (solid line) and ET (dashdot line).
@@
The distribution of the scalar mode with inclination angle, as shown in Fig. 8, exhibits an opposite trend to that of tensor and vector modes. It reaches its maximum value at $\iota=\pi/2$ and minimum values at $\iota=0,\pi$. The scalar mode outperforms the other two modes at $\iota=\pi/2$, while approaching zero at $\iota=0,\pi$. The patterns of variation in both tensor and vector modes are similar, although tensor modes show a greater amplitude of change. The SNR and FIM have been computed to determine the uncertainty of ppE parameters, as shown in Fig. 9.
@@
The scalar mode in Fig. 9 shows an opposite trend compared with other modes, as depicted in the distribution shown in Fig. 8. The scalar mode vanishes at $\iota=0,\pi$, resulting in a near absence of signal around these values. As a result, the parameter uncertainty for the scalar mode approaches infinity in Fig. 9. Additionally, considering the low-frequency limit, a $10^{5}\ \mathrm{M_{\odot}}$ MBHB with low-SNR demonstrates smaller $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ values compared with a $10^{7}\ \mathrm{M_{\odot}}$ MBHB with high-SNR. Owing to superior sensitivity and the lower transfer frequency, LISA exhibits significantly better $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ values than TianQin does. Ground-based detectors are unable to distinguish between the breathing mode and longitudinal mode; hence their corresponding $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ values are much larger than those of space-based detectors and become invalid.
@@
Besides the scalar mode, a higher SNR indicates a more robust signal with reduced parameter uncertainties. In our calculations, SNR is computed based on the GR case. Therefore, the SNR value is primarily influenced by tensor modes. In contrast to tensor and scalar modes, the uncertainty of the vector mode is significantly higher at $\iota=\pi/2$ for TianQin. The fixed orientation of TianQin limits variations in the response function at that specific angle, resulting in a significant increase in $\Delta\alpha_{V}$ near $\iota=\pi/2$.
@@
The transfer frequency not only directly impacts the degeneracy between breathing and longitudinal modes but also affects other aspects. The transfer function $\mathcal{T}$ is directly related to frequency, providing constraints beyond the low-frequency limit. $\alpha_{Q}$ and $\alpha_{D}$ determine the variation of the orbital angular frequency, and the influence of $\mathcal{T}$ can reduce parameter uncertainty. Therefore, the $\Delta\alpha_{Q}$ of $10^{5}\ \mathrm{M_{\odot}}$ MBHB is smaller than that for a $10^{6}\ \mathrm{M_{\odot}}$ MBHB, and $\Delta\alpha_{D}$ for both MBHBs also varies with the relationship near $\iota=\pi/2$. Because of the low-frequency limit, ground-based detectors are not affected by $\mathcal{T}$, causing the uncertainties $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ for SBBHs that depend solely on SNR. The aim of this section is to analyze the correlation between ppE parameter uncertainty and inclination angle for two detectors, one space-based and one ground-based.